### v1

# The Empty Chair: Using LLMs to Raise Missing Perspectives in Policy Deliberations

Abstract.
Deliberation is essential to well-functioning democracies, yet physical, economic, and social barriers often exclude certain groups, reducing representativeness and contributing to issues like group polarization. In this work, we explore the use of large language model (LLM) personas to introduce missing perspectives in policy deliberations. We develop and evaluate a tool that transcribes conversations in real-time and simulates input from relevant but absent stakeholders. We deploy this tool in a 19-person student citizens’ assembly on campus sustainability. Participants and facilitators found that the tool sparked new discussions and surfaced valuable perspectives they had not previously considered. However, they also noted that AI-generated responses were sometimes overly general. They raised concerns about overreliance on AI for perspective-taking. Our findings highlight both the promise and potential risks of using LLMs to raise missing points of view in group deliberation settings.
deliberation, policy, persona, language models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Computer Supported Cooperative Work; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Empirical studies in HCI
1. Introduction
Partisan gridlock and increasing political fragmentation have made effective policymaking increasingly difficult (Cohen, 2003; Jones, 2001). However, deliberative forums such as citizens’ assemblies have shown promise in bypassing party polarization and fostering productive discussions on contentious political issues (Caroline Bannock, 2019). Unfortunately, most deliberations do not take place in carefully structured settings with nationally representative participants. Instead, they often occur within homogeneous groups (McPherson et al., 2001). When this happens, deliberation can lead to group polarization, where individuals become more extreme in their initial positions rather than engaging with opposing viewpoints (Sunstein, 1999). This can be problematic if the goal of deliberation is to build common ground and consensus within a pluralistic electorate.
Given that large language models (LLMs) have demonstrated some fidelity in accurately responding to opinion surveys (Argyle et al., 2023; Park et al., 2024) and adopting different personas (Jiang et al., 2024), we explore whether an LLM-powered tool can help introduce missing perspectives in group deliberation. Specifically, we develop and test a system that transcribes discussions in real time and generates input from absent stakeholders. To evaluate its effectiveness, we deploy the tool in small group breakout sessions during a three-day citizens’ assembly with university undergraduates, examining whether it encourages participants to engage with perspectives that might otherwise be overlooked. From our results, we make several contributions for how to think about the design of computing technologies specifically in the context of small group policy deliberations.
  1. (1)
    LLM personas can help encourage participants to consider perspectives that may be missing from the conversation, making these viewpoints more concrete.
  2. (2)
    The way this technology is framed is crucial; presenting it as a tool to spark new discussions rather than as a replacement for or embodiment of a person or community is more likely to be effective.
  3. (3)
    There are risks in using AI to represent absent individuals or groups, and designers and users should be cautious not to treat these technologies as authentic representations of different perspectives.
2. Background
We cover the negative consequences of homogeneous deliberations, the role of perspective-taking in improving deliberations, and the opportunities and risks of using LLMs to predict opinions and model personas to encourage perspective-taking.
2.1. Group Polarization and Homogeneous Deliberation
While citizen deliberation holds promise as a means of effective policymaking and reducing political fragmentation, it can also reinforce polarization if certain conditions are not met (Sunstein, 1999). Group polarization—the tendency for individuals to adopt more extreme versions of their initial positions after discussion—has been well-documented across various contexts, including “attitudes, jury decisions, ethical reasoning, judgment formation, person perception, negotiation, and risk-taking” (Myers and Lamm, 1976). In a deliberative setting, this dynamic was observed during a “Deliberation Day” event in Colorado, where Sunstein et al. 2007 found that liberal participants became more left-leaning after discussion on three political issues, while conservative participants shifted further to the right. However, recent work suggests that with proper scaffolding, even homogeneous groups can depolarize (Khan and Siu, 2023). This indicates that, with the right structure and safeguards in place, group polarization in homogeneous deliberations is not inevitable.
2.2. Perspective-taking to Improve Deliberation
One promising approach to enhancing deliberation is perspective-taking, defined as the “active consideration of others’ mental states and subjective experiences”. Research has shown that perspective-taking can improve intergroup relations and reduce stereotype expression, among other benefits (Todd and Galinsky, 2014; Batson et al., 1997; Galinsky and Moskowitz, 2000). Computing systems have also been leveraged to foster perspective-taking. For instance, Khadpe et al. 2022 demonstrated that using a bot to regularly prompt team members to consider each other’s feelings during a collaboration exercise can enhance communication. Additionally, technology can expose individuals to different subjective experiences, fostering greater understanding and empathy. In the policy domain, Kim et al. 2019 crowd-sourced different perspectives on various policy issues and presented them in an engaging interface, enabling participants to interact with unexpected and differing viewpoints. Similarly, Yeo et al. 2024 improved deliberation quality by integrating reflection nudges into an online deliberation platform, finding that a persona-based approach was the most effective. This suggests that a persona-driven method could be particularly valuable in live, in-person deliberations, helping participants introduce and consider absent viewpoints in discussions.
2.3. LLMs as Personas and Representatives
Since our approach relies on personas to introduce missing perspectives, we briefly survey the literature on the benefits and limitations of using LLMs to model different groups and viewpoints. Several studies have demonstrated that prompted LLMs can reasonably predict survey responses from certain demographic and ideological groups (Argyle et al., 2023). However, they tend to perform worse when modeling minority perspectives (Santurkar et al., 2023) and often risk caricaturing these groups rather than authentically representing them (Cheng et al., 2023). Recent work suggests that providing LLMs with significantly more personal context improves their ability to predict individual survey responses and reduce this bias (Park et al., 2024). Additionally, LLMs have been shown to effectively generate text that aligns with prompted personas, adapting their style and content accordingly (Jiang et al., 2024). Given the demonstrated effectiveness of personas in prompting reflection to improve deliberation quality (Yeo et al., 2024) and the ability of LLMs to model different viewpoints with reasonable accuracy, we investigate whether a tool that dynamically generates personas and their responses in a deliberative setting can effectively help the group raise and understand missing perspectives.
3. Assembly Information
Since citizens’ assemblies are a key forum for policy deliberation, we deployed our tool in a relatively realistic setting: an in-person, three-day student assembly. The assembly convened 19 delegates to deliberate on the following remit:
[The university’s] commitment to achieving a zero-carbon campus by 2050, with an interim goal of net-zero emissions by 2026, underscores the urgency of implementing effective sustainability initiatives. What policies or initiatives should [the university] prioritize to enhance campus sustainability and meet its climate commitments? The assembly aims to generate actionable recommendations that position [the university] as a leader in campus sustainability and serve as a model for other institutions.
To maintain anonymity, we have redacted the university name and any identifying details about participants. The assembly followed a traditional citizens’ assembly structure: participants learned about the topic, heard from experts, deliberated, and ultimately formulated recommendations (Ferejohn, 2008). However, a key deviation from standard practice was the absence of a sortition process, which is typically used to ensure that participants are representative of the broader population (Delannoi and Dowlen, 2016). Without this process, selection bias likely influenced who chose to participate—favoring those with both the interest and the means to engage in such an event. Although delegates were compensated for their participation, those who signed up were already highly engaged with sustainability as an issue. As a result, our participant pool consisted of young, educated individuals with a preexisting commitment to environmental action. This introduced two notable limitations: (1) within the university, we lacked perspectives from staff, faculty, and other key stakeholders, and (2) at a broader level, several demographic groups—especially those less likely to prioritize sustainability—were underrepresented compared to the general U.S. population. However, this selection bias also created an ideal test case for our tool, which aimed to identify missing stakeholders and incorporate their perspectives into the discussion. By deploying our system in an assembly with a skewed participant pool, we could evaluate whether it effectively introduced absent viewpoints and broadened the scope of deliberation.
4. Tool Description
The practice of using an “empty chair” to represent an absent person or group is well-established in certain types of therapy (Harman, 1974) and is also a common technique in business settings, where user personas help teams understand different stakeholder needs (Nielsen, 2013). The idea to develop an AI-assisted empty chair tool for deliberative settings stemmed from observing a previous student assembly, where participants were tasked with role-playing predefined personas. However, they struggled to authentically embody the concerns and viewpoints of these personas. To address this challenge, we designed an AI-powered tool that dynamically generates and introduces missing perspectives into deliberation. Figure 1 outlines the process by which the tool creates dynamic personas that provide relevant input to the conversation. Using a transcript of the discussion along with brief contextual information about the topic and assembly setting, the tool identifies three stakeholder personas who may be affected by the assembly’s decisions. Each persona is displayed with demographic details—including race, gender, income, occupation, political leaning, and level of interest in sustainability (low, medium, high), as well as background story. We observed that common stakeholder personas included local small business owners and university facilities staff responsible for implementing proposed sustainability initiatives. Figure 2 provides an example of a generated persona. These personas are then integrated with the transcript to surface relevant points of disagreement and highlight missing perspectives within the discussion (see Figure 3 for an example). Additionally, since the tool was deployed during a session where participants were tasked with formulating questions for an expert panel, it also allowed users to generate a question from the perspective of a selected persona, helping to ensure that absent viewpoints were actively considered (see Figure 4 for an example).
Figure 1. The flow for generating missing stakeholders, stakeholder reflections, and finally stakeholder questions based on the assembly context and a live transcription of the conversation. The solid lines capture how the assembly context and the transcript helps generate missing stakeholder personas, which then is used to generate stakeholder reflections, and then finally the stakeholder questions.
Figure 2. An example persona generated from the conversation transcript and deliberation context. We have redacted any specific references to locations or universities.
Figure 3. An example of points of disagreement and missing perspectives from a persona.
Figure 4. An example question generated by a persona.
5. Tool Deployment
The tool was deployed during a one-hour session of the assembly, where participants were divided into four groups of 4–5 people. The session was structured into two main phases:
  1. (1)
    Initial Discussion and Question Drafting (30 minutes) – Participants discussed the expertise of the panel members and drafted potential questions to ask during the panel. During this time, the tool passively recorded and transcribed the conversation.
  2. (2)
    Tool Introduction and Stakeholder Engagement (30 minutes) – Facilitators provided a brief explanation of the tool’s purpose and functionality. They then used the tool to generate several stakeholder bios (see Figure 2). Delegates took turns reading each persona’s bio before selecting one to explore further.
Once a persona was selected, the facilitator clicked on it, revealing key points of disagreement and missing perspectives associated with that stakeholder (see Figure 3). Delegates read these perspectives and then engaged in a guided discussion about whether they should take this stakeholder’s concerns into account and, if so, how they might address them. Finally, the facilitator had the option to generate a question from the perspective of the selected stakeholder (see Figure 4). Delegates then discussed whether they should include this question in their list for the panel.
6. Evaluation
We evaluated the efficacy of the system via quantitative and qualitative feedback from student participants as well as interviews with the facilitators.
6.1. Activity Feedback from Delegates
To evaluate the tool’s effectiveness, we asked participants to complete a post-activity survey assessing their overall impressions and specific aspects of the tool (see Figure 5 for results). Overall, delegates found the activity engaging and thought-provoking ($m=5.83$, CI = [5.44, 6.23]) and considered it useful ($m=5.61$, CI = [5.19, 6.03]). They strongly agreed that the tool helped spark further discussion about missing perspectives ($m=6.16$, CI = [5.71, 6.60]) and found the system easy to use, with clear outputs ($m=6.05$, CI = [5.71, 6.39]). One of the most commonly cited benefits was the tool’s ability to introduce perspectives that might have otherwise been overlooked. As one student put it, “I liked the bringing in of more perspectives that we hadn’t thought of, since our assembly in particular is not representative.” Another student emphasized how the persona biographies made abstract concerns feel more tangible: “The persona’s background that shared the context behind their position, comments, and questions really helped to shape and represent a sort of real person that has a real problem.” Delegates particularly valued the points of disagreement and missing perspectives generated by the tool, with some describing it as a way to surface “blind spots on demand.” However, the AI-generated questions were seen as less useful ($m=5.42$, CI = [4.98, 5.87]), as some participants felt they were too general or redundant. One delegate explained, “Reading the person’s questions [was] least useful, as they seemed very general and sometimes already included in our questions.” Despite its strengths, the tool also raised concerns about the potential for AI to misrepresent perspectives or create a false sense of representation. One delegate warned that it “may have the potential to actually stifle diverse voices in citizen’s assemblies by acting as a poor substitute for actually bringing in diverse voices.” This delegate suggested that instead of framing the output as fully realized personas, the tool should “bring up missing points” to avoid the misleading impression that AI can perfectly represent all viewpoints, which was different than other delegates who appreciated the relatively detailed backgrounds of the personas. This tension reflects broader debates in social science research, where LLMs are increasingly used to model human perspectives (Hewitt et al., 2024). The risks are particularly pronounced given that LLMs tend to be less accurate at predicting the opinions of minority groups (Santurkar et al., 2023) and can introduce biases or caricatures depending on how personas are constructed (Gupta et al., 2024; Cheng et al., 2023). Addressing these challenges is critical, and we hope this work highlights both the potential benefits and the real risks of using AI to surface missing perspectives in deliberation.
6.2. Changes in Empathy and Understanding of Differing Opinions
Beyond gathering direct feedback on the tool, we sought to evaluate whether introducing missing perspectives through our system increased delegates’ empathy and understanding toward those with differing opinions. To assess this, we administered a pre-survey before the assembly to capture baseline attitudes and then provided the same survey shortly after the activity. As shown in Figure 6, delegates generally exhibited increased empathy toward individuals with different perspectives on sustainability following the activity ($m_{post}-m_{pre}=0.63$, $m_{post}=6.11\pm 0.42$, $m_{pre}=5.47\pm 0.51$). Their empathy also increased specifically toward those who outright disagreed with them ($m_{post}-m_{pre}=0.68$, $m_{post}=5.53\pm 0.48$, $m_{pre}=4.84\pm 0.50$). Additionally, participants were less likely to agree with the statement that individuals who do not prioritize sustainability are actively causing societal harm ($m_{post}-m_{pre}=-0.42$, $m_{post}=4.89\pm 0.49$, $m_{pre}=5.32\pm 0.56$). They also reported a greater ability to hear, understand, and respect arguments that conflicted with their own beliefs ($m_{post}-m_{pre}=0.63$, $m_{post}=6.00\pm 0.45$, $m_{pre}=5.37\pm 0.45$). However, because delegates engaged in other discussions throughout the day before using the empty chair tool, we cannot directly attribute these attitude shifts solely to tool usage. Qualitative feedback from participants indicated that changes in perspective were influenced both by interactions with other delegates and by engagement with the tool. In future work, we hope to measure this change in understanding of different positions in a randomized controlled experiment.
Figure 5. Survey responses capturing how participants felt about various aspects of the activity as well as their overall impressions
Figure 6. Changes in understanding towards different points between pre-survey (prior to first day) and post-survey (after first day). Red lines capture pre/post means, and blue lines and dots represent individuals.
6.3. Facilitator Interviews
In addition to participant feedback, we conducted semi-structured interviews with the four facilitators who introduced and guided the use of the tool. We asked them what worked well, what could be improved, and what they would like to see in future iterations. Overall, facilitators found the tool intuitive and efficient, describing it as “very easy to use … and very fast.” They appreciated how it helped surface overlooked perspectives, with one facilitator noting that it “quickly helps delegates go deeper into some perspectives that they totally missed.” Another highlighted its role in fostering more open discussions, explaining that “it got [the participants] to be more candid with one another.” One facilitator expressed enthusiasm for broader adoption, stating they “would like to see this activity being introduced more often.” Others suggested making the personas more interactive and allowing participants to engage with them throughout the assembly rather than limiting interactions to a single session. Facilitators also acknowledged the challenges of using AI to represent different people and perspectives, particularly the discomfort that can arise when striking a balance between realism and fiction. One facilitator proposed that framing the tool explicitly as a role-playing exercise might help mitigate these concerns, suggesting that a “design language that aligns with roleplay” would be beneficial since “roleplay is really normal.” They also emphasized the importance of setting clear expectations from the outset, noting that facilitators should clarify that “this is for you to empathize with other people.”
6.4. Influence on Question Generation
Since the primary goal of the session was to develop questions for the panel, we examined whether AI-generated suggestions were incorporated into the final list of recommended questions. Several personas represented local small business owners concerned about the economic impact of sustainability policies. These personas raised questions about affordability and the financial burden of sustainability initiatives, which delegates engaged with and ultimately recommended for the panel. For example, one group proposed the question: “Are there ways to implement sustainable practices in a way that are affordable for smaller institutions? Consider immediate economic implications on small business.” Another group asked: “How might the university’s environmental policies affect surrounding small businesses? How can we partner with local businesses for the initiative?” These were clear variations of the questions initially suggested by the AI-generated personas, such as: “How can the university support small businesses like mine in adopting sustainable practices without imposing additional costs?” “How can the university’s sustainability initiatives be designed to support and economically benefit small local businesses in surrounding communities, rather than imposing changes that may strain their budgets?” However, while these questions from the personas proved useful, a majority did not end up being selected to bring to the panel. Participants noted that some questions were too generic or redundant, suggesting that future iterations of the tool could improve the specificity of the questions it generates.
6.5. Participant Engagement with Personas
We also examined how participants reacted to the personas presented live during the deliberation. In several instances, delegates demonstrated perspective-taking by agreeing with, empathizing with, or acknowledging the value of the personas’ viewpoints. For example, in response to a persona representing a university facilities worker, one student noted: “The perspective of someone working in facilities that knows inefficiencies is a very valuable perspective to have because those people know more about where waste is occurring than, say, the president or these high-level researchers.” Similarly, a delegate engaged with a persona representing a small business owner (Tony), explaining: “It’s not that Tony doesn’t care about climate change at all. It’s just that he can’t afford to care about it in that regard.”. One participant even mentioned how they wanted to hear from a local business owner persona because it seemed to represent “people like my mom”. At times, participants disagreed with the personas, particularly when they perceived them as overly self-interested. Additionally, some personas were seen as less helpful when their perspectives were too similar to those already represented within the group. Both facilitators and participants expressed a preference for personas that introduced viewpoints significantly different from those of the delegates, as these were more likely to challenge assumptions and provoke meaningful discussion. By analyzing both the integration of AI-generated questions and the delegates’ responses to the personas, we gained insight into how the tool influenced deliberation. While it effectively introduced new perspectives and shaped the conversation, refining persona selection and improving question specificity could enhance its impact in future iterations.
7. Discussion
In this work, we explored the potential and risks of using AI-generated personas to introduce missing perspectives into small-group policy deliberations. Overall, participants responded positively, finding the tool useful for expanding their consideration of absent viewpoints and prompting discussions in new directions. Many appreciated how the personas made abstract concerns more concrete and surfaced overlooked perspectives, with facilitators noting that the tool encouraged deeper, more candid engagement. Delegates and facilitators both suggested that having the personas more dynamically respond to the conversation would be an interesting and useful addition to the current tool. However, two key concerns emerged. First, participants worried that AI-generated personas could misrepresent certain groups, potentially replacing rather than complementing authentic voices. Second, some found the personas’ contributions too generic or redundant, limiting their impact. Facilitators suggested that explicitly framing the tool either as a role-playing aid or as a mechanism for surfacing overlooked perspectives could help clarify its purpose. Future iterations could improve persona specificity and grounding by incorporating real-world data, such as community input or structured datasets, to enhance authenticity and relevance. While AI has the potential to enrich deliberation by introducing different viewpoints, careful design is needed to ensure it serves as a supplement rather than a substitute for genuine representation. We hope this work provides useful insights into the benefits and risks of deploying AI personas in a realistic deliberative settings and helps spark ideas for new applications of LLMS and personas in deliberations.
8. Limitations
There are several key limitations to our work. First, our study evaluation was primarily qualitative, with the primary goal being to elicit in-depth feedback in a relatively realistic setting (i.e. as part of a multi-day assembly with live, synchronous conversation). However, we did not run a controlled experiment, and thus future work should validate whether introducing AI personas is more effective than a baseline such as a role-playing activity. Second, our sample of university students may not be representative of how other populations may react to such a tool. Since the participants were relatively knowledgeable about the strengths and weaknesses of AI systems, their reactions may be quite different to those with less knowledge or familiarity with these systems. Finally, we did not solicit feedback from people whom the personas purport to represent (e.g. local business owners and university staff). Thus, although the points raised by the personas seemed plausible and effective at increasing perspective-taking, further validation would be needed to determine if they actually represent people or communities well.
9. Ethical Considerations
There are several risks associated with using AI personas to introduce missing perspectives in deliberation. First, if widely adopted, organizations may rely on AI as a substitute rather than an aid for recruiting participants with different views, potentially reducing efforts to ensure genuine representation. Additionally, AI personas are inherently imperfect and may misrepresent the people or groups they aim to portray. If their role is not clearly framed or their contributions are perceived as low quality, they could inadvertently undermine the perspectives they are meant to highlight, leading participants to view those viewpoints more negatively. Finally, if individuals whose perspectives are being represented have no agency or control over how they are depicted, this could further contribute to feelings of misrepresentation.
Acknowledgements.
Appendix A Technical Details
For all LLM responses, we used GPT-4o. Given that LLMs may have different capabilities in modeling various perspectives (Santurkar et al., 2023) we hope that in the future tools utilizing AI generated personas have various base model options that users can choose. Our app was deployed on an AWS EC2 instance that could be accessed from a public URL by facilitators.
Appendix B Prompts
Below we display the system and user prompts for generating stakeholders, stakeholder responses, and stakeholder questions.
B.1. Stakeholder Generation Prompt
System Prompt: You are an expert at identifying missing perspectives in discussions. Generate 3 detailed stakeholder bios for people whose perspectives are missing from the conversation. This is a conversation between undergraduate students. Each bio should include:
  • Name
  • Description
  • Demographics:
    • Age
    • Gender
    • Income
    • Education
    • Profession
    • Political Leaning (Left, Right, Center)
    • Sustainability Interest (Low, Medium, High)
The response should be formatted as a JSON object with a ‘‘stakeholders’’ array containing objects with ‘‘name’’, ‘‘description’’, and ‘‘demographics’’ fields.
In addition to their thoughts and experience on sustainability, the description should also include personal details about the stakeholder, specifically:
  1. (1)
    Their background and experiences
  2. (2)
    Their beliefs and values
  3. (3)
    Their personal experiences
  4. (4)
    Their thoughts on sustainability
  5. (5)
    Their thoughts on the university’s sustainability policies
Be fairly detailed in the description, including their background, experiences, and beliefs. The demographics should be a JSON object with the specified fields. Make sure to include people who may not be interested or believe in sustainability.
User Prompt: Based on this transcript, generate 3 missing stakeholders. The discussion, taking place among university undergraduate students, has the following theme:
{THEME}
The stakeholders should be individuals who may have a vested interest in the university’s sustainability policies but are not undergraduate students. For example, they may be other types of university affiliates, people in the Boston area, or individuals who are not affiliated with the university but are still impacted by sustainability policies. At least one stakeholder should have low sustainability interest. Format the response as a JSON object with a ’stakeholders’ array containing objects with ‘‘name’’, ‘‘description’’, and ‘‘demographics’’ fields. The conversation is as follows:
{transcript}
B.2. Stakeholder Reflections Prompt
System Prompt: You are an expert at writing authentic reflections from different perspectives, considering their unique backgrounds, beliefs, and experiences.
User Prompt: You are an expert at understanding and expressing different perspectives. Given a stakeholder’s profile and a conversation transcript, generate a detailed reflection from their perspective. The reflection should feel authentic to their background and personality. The conversation is part of a broader assembly, the theme of which is discussing sustainability policy at the university.
Stakeholder Profile:
  • Name: {stakeholder[‘‘name’’]}
  • Description: {stakeholder[‘‘description’’]}
  • Demographics:
        {json.dumps(stakeholder["demographics"], indent=2)\}
The conversation transcript is as follows:
{transcript}
Generate a reflection from this stakeholder’s perspective that includes:
  1. (1)
    What they agree with and why (based on their background and experiences)
  2. (2)
    What they disagree with and why (based on their background and experiences)
  3. (3)
    What they think is missing from the conversation
The reflection should be written in the first person and should authentically reflect the stakeholder’s:
  • Education level and profession
  • Political views ({stakeholder[demographics][political_leaning]})
  • Interest in sustainability ({stakeholder[demographics][sustainability_interest]})
  • Personal background and experiences
  • Socioeconomic status
Return the reflection as a JSON object formatted with the following keys:
  • ‘‘agree_explanation’’: Explanation of what they agree with (about 150 words)
  • ‘‘disagree_explanation’’: Explanation of what they disagree with (about 150 words)
  • ‘‘missing_perspectives’’: Explanation of what they think is missing from the conversation (about 150 words)
Each explanation should be specific to what was discussed in the conversation. Do not be overly polite or positive. Do not mention demographics explicitly in the reflection.
Appendix C Generating Stakeholder Questions
System Prompt: You are an expert at generating insightful questions from different perspectives.Given a stakeholder’s profile and a conversation transcript, generate a question that this stakeholder would likely ask, based on their background, demographics, interests, and their reflection on the discussion. The question should be relevant to the topic of university sustainability and should add a new perspective to the conversation. The question should be directed at an expert if possible. Extrapolate the stakeholder’s perspective from the transcript and demographics and ask questions someone like this would ask, but don’t be obvious about it. Keep the questions relatively general, not too specific to the background of the stakeholder. Keep the questions relatively short and concise.
User Prompt:
Stakeholder Profile:
  1. (1)
    Name: stakeholder[“name”]
  2. (2)
    Description: stakeholder[“description”]
  3. (3)
    Demographics: json.dumps(stakeholder[“demographics”], indent=2)
{reflection_text}
Current Conversation: {transcript}
Current Questions: {current_questions_text}
The experts are: {EXPERTS}
Generate a question that this stakeholder would ask, based on their perspective, background, and reflection on the discussion. The question should align with their points of disagreement and what they think is missing from the conversation. Format the response as a JSON object with “question”, “explanation”, and “expert” fields, where expert is the expert the question should be directed at. The explanation should justify why this stakeholder would ask this question